OpenStudy (anonymous):

according to this property, 5a + a =5a +1a. what is the property?

OpenStudy (anonymous):

communative property

OpenStudy (anonymous):

Commutative*

OpenStudy (anonymous):

that would be the identity property of 1 under multiplication

OpenStudy (anonymous):

commutative would say \[5a+a=a+5a\]

OpenStudy (anonymous):

derp

OpenStudy (anonymous):

i googled it and got communative

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Commutative_property Identity property http://www.learningwave.com/lwonline/numbers/identity.html not sure what it is but I would say Identity property

OpenStudy (anonymous):

k then go to www.purple math.com

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still commutative http://www.purplemath.com/modules/numbprop.htm

OpenStudy (anonymous):

thats what i said google doesn't lie :)

OpenStudy (anonymous):

google doesn't lie, but it doesn't tell the truth either. it only gives you results based on what you type in

OpenStudy (anonymous):

the commutative property of addition says \[a+b=b+a\] order doesn't matter

OpenStudy (anonymous):

for multiplications it says \[ab=ba\]

OpenStudy (anonymous):

if you look at the difference between the expression \[5a+a\] and \[5a+1a\]the order is the same. the only difference is the introduction of the 1

OpenStudy (anonymous):

that is of addition im talking communative property.

OpenStudy (anonymous):

If you type in communative property it says oh you must have mis-typed commutative let's show you results for commutative property

OpenStudy (anonymous):

communative property doesn't exist mate only commutative

OpenStudy (anonymous):

ok i'm causing an augument BYE

OpenStudy (anonymous):

so the only question is "why is \[a=1a\] ?" and the answer is because \[1a=a1=a\] that is the identity property of 1 under multiplication

OpenStudy (anonymous):

i went with commutative...

OpenStudy (anonymous):

well go with what you like, but it is certainly wrong

OpenStudy (anonymous):

I would strongly suggest identity